OpenStudy (anonymous):

How much money invested at 5% compounded continuously for 3 years will yield $820? Answer a. $952.70 b. $818.84 c. $780.01 d. $705.78

OpenStudy (nurali):

$820=Pe^.15 $820=P*1.16 P=820/1.16=$706.90

OpenStudy (kropot72):

\[A=Pe ^{rt}\] \[P=\frac{820}{e ^{0.05\times 3}}=705.78\]